OpenStudy (anonymous):

Giving that tan 4pi/9=5.6713, determine the following cot pi/18 and tan 13pi/9

OpenStudy (mertsj):

\[\frac{13\pi}{9}=\frac{4\pi}{9}+\pi \]

OpenStudy (mertsj):

Do you mean cot 4pi/18?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

Giving that tan 4pi/9=5.6713, determine the following a) cot pi/18 and b ) tan 13pi/9

OpenStudy (anonymous):

answer should be a) 5.67; cot pi/18 = cot(pi/2-4pi/9)= tan 4pi/9 b) 5.67; tan 13pi/9= tan (3pi/2-pi/18) = cot pi/18